# Also at ]INFN - Sezione di Milano Also at ]INFN - Sezione LNGS A volcanic chronosequence as a time-resolved paleo-detector array to study the cosmic-ray flux in the Late Pleistocene and Holocene

### v2

@@ Abstract
We present a phenomenological study demonstrating the feasibility of using olivine xenoliths from the Chaîne des Puys as a time-resolved paleo-detector array to probe the cosmic-ray flux over the last 40,000 years. This volcanic region provides a unique chronosequence of samples brought to the surface by well-dated eruptions. By modeling the expected density of nuclear recoil tracks induced by cosmic-ray muons in olivine, we show that the signal is detectable and above backgrounds from natural radioactivity. We demonstrate that by analyzing samples with different exposure ages, it is possible to construct a time-differential measurement of the cosmic-ray flux. This method shows sensitivity to historical variations, such as the enhanced flux expected during the Laschamp geomagnetic excursion (41 kyr) and the potential contribution from nearby supernovae, for which we use the Antlia supernova remnant precursor as a benchmark. This work establishes a new application of the paleo-detector technique for long-scale time-domain high-energy astrophysics and provides the direct scientific motivation for experimental efforts to measure these track records.
I Introduction
The emerging field of paleo-detectors aims to utilize natural minerals as long-exposure particle detectors, offering a unique opportunity to probe astrophysical phenomena and fundamental particle physics over geological timescales [Baum2023]. The concept underlying this field is the search for linear defects left by particles in the crystalline structure of suitable minerals, called paleo-detectors. This idea has been proposed in recent years, as a technique to detect dark matter [Baum:2018tfw, Drukier:2018pdy] and neutrinos [Baum:2019fqm]. These studies take advantage of the enormous exposure that can be acquired through age, even with a small amount of material.
The primary mechanism of interest for cosmic-ray studies using paleo-detectors is the interaction of high-energy secondary particles from atmospheric showers, predominantly muons, with the nuclei of the crystal lattice. While a muon passage itself does not leave a track, its interaction can induce spallation, causing the target nucleus or a fragment of it to recoil with significant kinetic energy. This recoiling nucleus, often a new cosmogenic isotope, travels a short distance through the lattice, leaving behind a damage track with a length ranging from hundreds of nanometers to tens of micrometers. The density of these accumulated tracks is proportional to the integrated particle flux, while the distribution of their lengths is correlated with the recoil energy spectrum of the produced isotopes. By measuring these features, one can, in principle, reconstruct the history of the cosmic-ray flux at Earth, providing a novel tool for probing high-energy astrophysical events [Galelli:2025, Caccianiga2024].
@@ I Introduction
In this paper, we extend and evolve this methodology from a single-epoch measurement to a time-resolved analysis. We present a new phenomenological study focused on olivine xenoliths from the Chaîne des Puys volcanic field in Auvergne, France. This region provides a unique geological scenario: a chronosequence of samples brought to the surface by a series of well-dated, distinct volcanic eruptions spanning the late Pleistocene and early Holocene (40,000 to 7,000 years ago) [Miallier2012, Boivin2017]. During eruptions, the host basaltic magma, with temperatures exceeding 1000°C, acts as a natural furnace, annealing any pre-existing tracks from the xenolith’s previous geological history. This process ensures that the track record in any given sample begins precisely at the time of its host eruption. Subsequent eruptions in the volcanic chain source mantle material of similar characteristics, creating a series of well-defined, sequential, and overlapping exposure windows. We aim to establish the experimental feasibility of this approach and lay the groundwork for transforming the paleo-detector concept into a time-resolved observatory for probing the recent history of our galactic neighborhood.
II The Geological Scenario: Volcanism in the Chaîne des Puys
@@
| Volcanic Event | Age (kyr) | Reference |
@@
III Modeling Track Production in Olivine
To quantify the expected signal in the Chaîne des Puys xenoliths, we have developed a comprehensive simulation pipeline based on the framework established in our previous work [Caccianiga2024]. The full analysis code is publicly available111The complete analysis code and input files for this work can be found at https://github.com/cgalelli/PrimusCode.. The goal of the pipeline is to translate a given cosmic-ray flux history and geological scenario into the observable number and length distribution of nuclear recoil tracks per unit mass.
The simulation chain is a three-step process. First, we model the primary cosmic-ray flux and its propagation through the atmosphere. This is handled by the MCEq (Matrix Cascade Equation) package [Fedynitch2015, Riehn:2017], which calculates the resulting sea-level muon flux. The primary cosmic-ray spectrum input to MCEq is controlled by the crflux package 222https://crfluxmodels.readthedocs.io/en/latest/. To simulate historical variations, we have implemented a custom class inheriting from crflux.PrimaryFlux. This allows us to apply on-the-fly modifications to a baseline model, enabling the simulation of scenarios such as an enhanced flux from a nearby supernova or from a geomagnetic excursion like the Laschamp event. For this work, we simulate a ”normal” scenario consistent with the currently measured cosmic-ray spectrum using crflux.HillasGaisser2012 [Gaisser2012], and an alternative scenario (”SN250”) including an additional, now-extinct cosmic-ray component from a supernova with parameters consistent, as a possible realistic example, with the Antlia supernova: age 50 kyr, distance 250 pc [Fesen2021].
@@ IV Results and discussion
The ultimate goal of this method is to probe the time-variability of the cosmic-ray flux. In figure 5, we plot the integrated track number as a function of the exposure time for each volcanic scenario. As expected, the track count scales linearly with exposure duration, with slight deviations due to time integration steps of the astrophysical scenarios. The blue and dotted lines represent the expected track accumulation for the normal and SN250 flux models, respectively. The two scenarios are discriminated at least at the 1 level even when accounting for a possible 30% counting systematic. For this predicted enhancement for samples exposed during the Laschamp event, modeled by an additional primary flux component in the GeV to 10 GeV range. This additional enhancement lands inside the 1 confidence interval even without accounting for systematics, and as such, should be hard to detect with this method.
V Conclusions
@@
This work also shows that the chronosequence of eruptions allows for a differential measurement, enabling a direct probe of the flux’s time-variability during the Late Pleistocene and Holocene. This technique shows possible sensitivity to specific phenomena, such as the enhanced flux expected during the Laschamp geomagnetic excursion and the potential contribution from nearby, now-extinct sources like the Antlia supernova.
This phenomenological work provides the direct scientific antecedent for the PRImuS (Paleo-astroparticles Reconstructed with the Interactions of MUons in Stone) project [Galelli:2025], an experimental effort funded by the INFN. The primary goal of PRImuS is to realize the measurements proposed in this paper and in our previous work on Halite. To this end, we have already begun the crucial first step of sample recovery, starting with olivine xenoliths sourced directly from the 7,650-year-old Puy de Montcineyre maar. The successful identification of muon-induced tracks in these and other samples from the Chaîne des Puys will mark a significant advance, transforming the paleo-detector concept into a new observational tool for time-domain high-energy astrophysics.

### v3

# A volcanic chronosequence as a time-resolved paleo-detector array to study the cosmic-ray flux in the Late Pleistocene and Holocene

@@ Abstract
We present a phenomenological study demonstrating the feasibility of using olivine xenoliths from the Chaîne des Puys as a time-resolved paleo-detector array to probe the cosmic-ray flux over the last 40,000 years. This volcanic region provides a unique chronosequence of samples brought to the surface by well-dated eruptions. By modeling the expected density of nuclear recoil tracks induced by cosmic-ray muons in olivine, we show that the signal is detectable and above backgrounds from natural radioactivity. We demonstrate that by analyzing samples with different exposure ages, it is possible to construct a time-differential measurement of the cosmic-ray flux. This method shows sensitivity to historical variations, such as the enhanced flux expected during the Laschamp geomagnetic excursion (41 kyr) and the potential contribution from nearby supernovae, for which we use the Antlia supernova remnant precursor as a benchmark. This work establishes a new application of the paleo-detector technique for long-scale time-domain high-energy astrophysics and provides direct scientific motivation for experimental efforts to measure these track records.
1 Introduction
The emerging field of paleo-detectors aims to utilize natural minerals as long-exposure particle detectors, offering a unique opportunity to probe astrophysical phenomena and fundamental particle physics over geological timescales [Baum2023]. The concept underlying this field is the search for linear defects left by particles in the crystalline structure of suitable minerals, called paleo-detectors. This idea has been proposed in recent years, primarily as a technique to detect dark matter [Baum:2018tfw, Drukier:2018pdy] and neutrinos [Baum:2019fqm]. These studies take advantage of the enormous exposure that can be acquired through age, even with a small amount of material.
The primary mechanism of interest for cosmic-ray studies using paleo-detectors is the interaction of high-energy secondary particles from atmospheric showers, predominantly muons, with the nuclei of the crystal lattice. While a muon passage itself does not leave a track, its interaction can induce spallation, causing the target nucleus or a fragment of it to recoil with significant kinetic energy. This recoiling nucleus, often a new cosmogenic isotope, travels a short distance through the lattice, leaving behind a damage track with a length ranging from hundreds of nanometers to tens of micrometers. The density of these accumulated tracks is proportional to the integrated particle flux, while the distribution of their lengths is correlated with the recoil energy spectrum of the produced isotopes. By measuring these features, one can, in principle, reconstruct the history of the cosmic-ray flux at Earth, providing a novel tool for probing past high-energy astrophysical events [Galelli:2025, Caccianiga2024].
@@ 1 Introduction
In this paper, we extend and evolve this methodology from a single-epoch measurement to a time-resolved analysis. We present a new phenomenological study focused on olivine xenoliths, inclusions from mantle minerals in lavic flows, from the Chaîne des Puys volcanic field in Auvergne, France. This region provides a unique geological scenario: a chronosequence of samples brought to the surface by a series of well-dated, distinct volcanic eruptions spanning the late Pleistocene and early Holocene (40,000 to 7000 years ago) [Miallier2012, Boivin2017]. During eruptions, the host basaltic magma, with temperatures exceeding 1000°C, acts as a natural furnace, annealing any pre-existing tracks from the xenolith’s previous geological history. This process ensures that the track record in any given sample begins precisely at the time of its host eruption. Subsequent eruptions in the volcanic chain source mantle material of similar characteristics, creating a series of well-defined, sequential, and overlapping exposure windows. We aim to establish the experimental feasibility of this approach and lay the groundwork for transforming the paleo-detector concept into a time-resolved observatory for probing the recent history of our galactic neighborhood.
2 The Geological Scenario: Volcanism in the Chaîne des Puys
@@
| Volcanic Event | Age (kyr) | References |
@@
3 Modeling Track Production in Olivine
To quantify the expected signal in the Chaîne des Puys xenoliths, we have developed a comprehensive simulation pipeline based on the framework established in our previous work [Caccianiga2024]. The full analysis code is publicly available222The complete analysis code and input files for this work can be found at https://github.com/cgalelli/PrimusCode.. The goal of the pipeline is to translate a given cosmic-ray flux history and geological scenario into the observable number and length distribution of nuclear recoil tracks per unit mass.
The simulation chain is a three-step process. First, we model the primary cosmic-ray flux and its propagation through the atmosphere. This is handled by the MCEq (Matrix Cascade Equation) package [Fedynitch2015, Riehn:2017], which calculates the resulting sea-level muon flux. The primary cosmic-ray spectrum input to MCEq is controlled by the crflux package 333https://crfluxmodels.readthedocs.io/en/latest/. To simulate historical variations, we have implemented a custom class inheriting from crflux.PrimaryFlux. This allows us to apply on-the-fly modifications to a baseline model, enabling the simulation of scenarios such as an enhanced flux from a nearby supernova or from a geomagnetic excursion like the Laschamp event. For this work, we simulate a ”normal” scenario consistent with the currently measured cosmic-ray spectrum using crflux.HillasGaisser2012 [Gaisser2012], and an alternative scenario (”SN250”) including an additional, now-extinct cosmic-ray component from a supernova with parameters consistent, as a possible realistic example, with the Antlia supernova: age 50 kyr, distance 250 pc [Fesen2021].
@@ 4 Results and discussion
The ultimate goal of this method is to probe the time-variability of the cosmic-ray flux. In figure 5, we plot the integrated track number as a function of the exposure time for each volcanic scenario. As expected, the track count scales linearly with exposure duration, with slight deviations due to time integration steps of the astrophysical scenarios. The blue and dotted lines represent the expected track accumulation for the normal and SN250 flux models, respectively. The two scenarios are discriminated at least at the 1 level even when accounting for a possible 30% counting systematic. For the predicted enhancement for samples exposed during the Laschamp event, modeled by an additional primary flux component in the GeV to 10 GeV range, the signal difference lands close to the 1 confidence interval even without accounting for systematics, and as such, should be hard to detect with this method.
5 Conclusions
@@
This work also shows that the chronosequence of eruptions allows for a differential measurement, enabling a direct probe of the flux’s time-variability during the Late Pleistocene and Holocene. This technique shows possible sensitivity to specific phenomena, such as the potential contribution from nearby, now-extinct sources like the Antlia supernova, and in more optimistic cases, to weaker phenomena like the enhanced flux expected during the Laschamp geomagnetic excursion.
This provides a direct scientific antecedent for the PRImuS (Paleo-astroparticles Reconstructed with the Interactions of MUons in Stone) project [Galelli:2025], an experimental effort funded by the INFN. The primary goal of PRImuS is to realize the measurements proposed in this paper and in our previous work on Halite. To this end, we have already begun the crucial first step of sample recovery, starting with olivine xenoliths sourced directly from the 7,650-year-old Puy de Montcineyre maar. The successful identification of muon-induced tracks in these and other samples from the Chaîne des Puys will mark a significant advance, transforming the paleo-detector concept into a new observational tool for time-domain high-energy astrophysics.
Acknowledgments
PRImuS is an INFN experiment funded by the CSN5 Young Scientist Grant 2024. The authors thank Didier Miallier, Thierry Pilleyre, Pierre-Jean Gauthier, Valentin Niess, Denis Andrault, and Emmanuel Gardes for the fruitful discussions in Clermont-Ferrand on xenoliths and olivine samples from the Chaîne des Puys and possible methods of analysis and readout. We also thank the members of the paleo-detector community, especially Alexey Elykov and Emilie LaVoie-Ingram, for their interest and exchange of ideas.